# Understanding the Issues, Their Causes and Solutions in Microservices Systems: An Empirical Study

### v2

@@ Abstract
Index Terms:
Microservices System, Microservices Architecture, Issues, Open Source Software, Empirical Study
@@ 1 Introduction
Motivating Example: We now contextualize the issues, causes, and solutions based on an example illustrated in Figure 1. The example is taken from the Spinnaker project, an open-source microservices project hosted on GitHub (see Table I), and annotated with numbering to represent a sequence among the reported issue, its cause(s), and the solution(s) to resolve the issue. Figure 1 shows auxiliary information about the Spinnaker project, such as project description, stars, and contributors. As shown in the example, a contributor, typically a microservices developer, writes code and may provide additional details of the code through comments. Once the code is compiled, the contributor reports permission denied issue highlighting “pipeline save with Admin account fails with permission denied”. As the next step, the same or other contributor highlights the cause for such issue as “Spinnaker user does not have access to the service account”. As the last step, an individual or a community of developers provides a solution such as “Allow the admin users to save the accounts” to resolve the issue. We are only interested in analyzing issues that have been marked as closed to ensure that a solution to resolve the issue exists. As shown in Table I, the Spinnaker project has 4,595 closed issues and 121 contributors.
@@ 2.1 Research Questions
{tcolorbox}
[sharp corners, boxrule=0.1mm,] RQ1: What issues do occur in the development of microservice systems?
@@
{tcolorbox}
[sharp corners, boxrule=0.1mm,] RQ2: What are the causes of issues that occur in microservices systems? Rationale: The aim of RQ2 is to investigate and classify the root causes behind the issues identified in RQ1 and map causes to issues. The answer to RQ2 helps practitioners avoid common issues in microservices systems.
{tcolorbox}
[sharp corners, boxrule=0.1mm,] RQ3: What solutions are proposed to fix issues that occur in microservices systems? Rationale: The aim of RQ3 is to identify the solutions for the issues according to their causes and to develop the taxonomy of solutions. The answer to RQ3 helps to understand the fixing strategies for addressing microservices issues.
{tcolorbox}
[sharp corners, boxrule=0.1mm,] RQ4: What are the practitioners’ perspectives on the taxonomies of the identified issues, causes, and solutions in microservices systems? Rationale: The taxonomies of issues, causes, and solutions developed in RQ1, RQ2, and RQ3 are based on 15 open-source microservices systems and interviewing 15 practitioners. RQ4 aims to evaluate the taxonomies of issues, causes, and solutions built in RQ1, RQ2, and RQ3 by conducting a relatively large-scale online survey.
@@ 2.2.1 Step A – Identify & Select MSA-based OSS Projects
We contacted a total of 426 contributors, with 39 of them responded to our query. Based on their confirmation, we selected 15 MSA-based OSS projects (see Table I). These 15 selected projects, based on the available data at the time of project selection, represent the highest number of issues, making them highly relevant to our study.
TABLE I: Identified open-source microservices systems
@@
Spinnaker
4595 121 1.2K 8.5K
Cortex
1120 226 681 4.7K
Jaeger
995 239 1.9K 15.8K
eShopOnContainers
986 157 8.9K 20.7K
Goa
930 97 485 4.7K
Light-4j
584 37 588 3.4K
Moleculer
473 102 497 5.1K
Microservices-demo
287 55 2.1K 3.1K
Cliquet
207 19 20 65
Deep-framework
174 12 75 537
Scalecube
130 20 90 547
Lelylan
123 7 93 1.5K
Open-loyalty
175 14 80 300
Spring PetClinic
69 32 1.4K 1.1K
Pitstop
39 15 490 890
@@ 2.2.2 Step B – Synthesize Issues, Causes, and Solutions
Raw Data Collection: We chose 15 microservices projects (see Table I) as the source for building the dataset to answer our RQs. These 15 projects were chosen because they are significantly larger than other microservices projects hosted on GitHub. Hence, it is highly likely that their contributors had more discussions about the type of issues, causes, and solutions in issue tracking systems. The discussions relating to a software system can usually be captured in issue tracking systems [24]. We initially extracted 10,222 issue titles, issue links, issue opening and closing dates, and the number of contributors for each issue through our customized Python script (see the Raw Data sheet in [19]). We stored this information in MySQL and exported it into MS Excel sheets for further processing. We extracted only closed issues because it could increase the chances of answering all our RQs (e.g., solutions).
@@ 2.3.3 Data Analysis
TABLE II: Overview of the interviewees and their demographic information
@@
P1
Software Architect, Developer
Python, Java, Node.JS
E-commerce, Healthcare
14 Years
6 Years
Sweden
P2
Developer
Java with Spark
E-commerce
9 Years
5 Years
USA
P3
Software Architect
Python, Go, Java
E-commerce
13 Years
6 Years
UK
P4
Software Architect, Developer
Python, Java
Educational ERP
7 Years
3 Years
Pakistan
P5
Software Architect
Python, Java
Internet of Things
12 Years
7 Years
Canada
P6
Software Architect, Developer
Java, Node.JS, Python
Healthcare
15 Years
8 Years
Australia
P7
Software Engineer
C#.Net
E-commerce, Banking
10 Years
4 Years
Canada
P8
DevOps Consultant
Kotlin, Python
Network applications
9 Years
6 Years
Canada
P9
Software Architect
Java
Education, Healthcare
5 Years
4 Years
Chile
P10
Application Developer, Architect
Java
Financial (Insurance)
10 Years
8 Years
Sweden
P11
DevOps Consultant
Java, Kotlin,Python
Telecommunication
6 Years
3 Years
Norway
P12
Application Developer
Angular, Puthon
Manufacturing
8 Years
4 Years
China
P13
Principal Consultant
Swift
Transportation
9 Years
5 years
USA
P14
Azure Technical Engineer
JavaScript, Golang
Embedded systems
12 Years
4 Years
Pakistan
P15
Software Architect
Ruby, UML
Payment applications
16 Years
7 Years
India
@@ 3 Results
This section presents the analyzed results of this study, addressing the four RQs outlined in Section 2.1. The analyzed results are further organized as categories (e.g., Technical Debt), subcategories (e.g., Code Debt), and types (e.g., code smell). We present categories in boldface, subcategories in italic, and types in small capitals. The relevant examples are provided as quoted messages along with their issue ID numbers to facilitate the traceability to our dataset (see the Initial Codes sheet in [19]). We report the types of issues in Section 3.1 (see Figure 4), the types of causes in Section 3.2 (see Table III), the types of solutions in Section 3.3 (see Table IV), and practitioners’ perspective on these taxonomies in Section 3.4 (see Table V).
@@ 3.1 Types of Issues (RQ1)
{tcolorbox}
[sharp corners, boxrule=0.1mm,] Key Findings of RQ1: We identified 2,641 instances of issues by mining developer discussions in 15 open-source microservices systems with 48 instances of issues mentioned by the interviewees and 9 instances of issues mentioned by the survey participants, which are 2,698 issues in total. The issue taxonomy consists of 19 categories, 54 subcategories, and 402 types, indicating the diversity of the issues in microservices systems. The majority of issues are related to Technical Debt (25.46%), CI/CD (11.60%), and Exception Handling (8.45%).
Figure 4: A taxonomy of issues in microservices systems
@@ 3.2 Causes of Issues (RQ2)
The taxonomy of causes of microservices issues is provided in Table III. It is worth mentioning that not all the issue discussions provide the information about their causes. Therefore, we identified 2,225 issue discussions containing information about the causes. The taxonomy of causes is derived by mining developer discussions (i.e., 2,225 instances of causes), conducting practitioner interviews (i.e., 31 instances of causes, see Section 2.3.3), and conducting a survey (i.e., 11 instances of causes, see Section 3.4). Hence, we got a total of 2,267 instances of causes. We identified a total of 228 types of causes that can be classified into 8 categories and 26 subcategories. Due to space limitations, we only list the top two types of causes for each subcategory in Table III. The detail of the types of causes can be found in the dataset [19]. The results show that General Programming Error (860 out of 2267), Missing Features and Artifacts (386 out of 2267), and Invalid Configuration & Communication Problems (382 out of 2267) are the top three categories of causes. Each cause category is briefly discussed below.
1. General Programming Error (GPE) (860/2267, 37.93%): This category captures the causes that are based on a broad range of errors occurred in different phases of microservices system development, such as coding (e.g., syntax errors), testing (e.g., incorrect test cases), and maintenance (e.g., wrong examples in documentation) phases. We identified and classified 78 types of GPE causes in six subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
2. Missing Features and Artifacts (MFA) (386/2267, 17.02%): This category represents the causes behind microservices systems issues that occur due to missing required features, packages, files, variables, and documentation and tool support. We identified and classified 27 types of MFA causes in four subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 29 types of causes in two subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 28 types of causes in five subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 17 types of SD&IA causes in 2 subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 21 types of PSM causes in 3 subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
7. Insufficient Resources (IR) (93/2267, 4.10% ): Microservices systems are at risk of delivering the required outcome without sufficient resources. We identified and classified 4 types of IR causes in two subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
8. Fragile Code (FC) (24/2267, 1.05%): Fragile code refers to code that is difficult to change, and a minor modification in fragile code may break the service or module. We identified and classified 6 types of FC causes in two subcategories (see Table III and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
TABLE III: Taxonomy of causes of issues in microservices systems
@@
{tcolorbox}
[sharp corners, boxrule=0.1mm,] Key Findings of RQ2: We found that not all the issue discussions provide the information about their causes, and finally we identified 2,225 issue discussions containing information about the causes with 31 causes mentioned by the interviewees and 11 causes indicated by the survey participants, which are 2,267 causes instances in total. The cause taxonomy of microservices issues consists of 8 categories, 26 subcategories, and 228 types. The majority of causes are related to General Programming Errors (37.93%), Missing Features and Artifacts (17.02%), and Invalid Configuration and Communication (16.85%).
@@ 3.3 Solutions of Issues (RQ3)
The taxonomy of solutions for microservices issues is provided in Table IV. It is worth mentioning that not all the issue discussions provide the information about their solutions. Therefore, we identified 1,899 issue discussions containing the information about the solutions. The taxonomy of solutions is derived by mining developer discussions (i.e., 1,899 solutions), conducting practitioner interviews (i.e., 36 solutions, see Section 2.2.2), and conducting a survey (i.e., one instance of solution, see Section 3.4). Hence, we got a total of 1,936 solutions. We identified a total of 196 types of solutions that can be classified in 8 categories and 35 subcategories. Due to space limitations, we only list the top two types of solutions for each subcategory in Table IV. The detail of the types of solutions can be found in the dataset [19]. The results show that Fix Artifacts (1056 out of 1936), Add Artifacts (360 out of 1936), and Modify Artifacts (210 out of 1936) are the top three categories of solutions. Each solution category is briefly discussed below.
1. Fix Artifacts (1056/1936, 54.54%): During the analysis of developer discussions about microservices issues, we identified that most developers did not explicitly mention any solution to the problems. They fixed the issue in the local repository and send the fixed code (e.g., through a pull request) to the maintainer of the public repository. In this case, from the developer discussions, we could not find exactly what they added, removed, or modified in the project to fix a specific issue. Therefore, we named this category Fix Artifacts. We identified and classified 25 types of fix artifacts solutions in 4 subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 33 types of solutions in 10 subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
3. Modify Artifacts (212/1936, 10.95%): Besides adding new artifacts to the existing system to address the microservices issues, we also identified a large number of solutions in which the developers explicitly mentioned how they modified modules, services, packages, APIs, scripts, methods, objects, data types, identifiers, databases, and documentation to address the microservices issues. We identified and classified 31 types of solutions in 5 subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
4. Remove Artifacts (39/1936, 2.01%): This category covers the solutions in which artifacts are removed to address several types of microservices issues. We identified and classified 14 types of solutions in 5 subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
5. Manage Infrastructure (160/1936, 8.26%): This category captures the solutions that are based on efficient resource utilization for addressing the microservices issues. We identified and classified 20 types of solutions in 2 subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
6. Manage Configuration and Execution (50/1936, 2.58%): Managing configuration and execution enables developers to track changes in microservices systems and their consuming applications over time, for example, the ability to track the version history of configuration changes for multiple instances of microservices systems. We identified and classified 13 types of solutions in 2 subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 21 types of solutions in two subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
8. Import/Export Artifacts (12/1936, 0.61%): We found several issues that can be fixed by importing and exporting various artifacts. We identified and classified 3 types of solutions in 2 subcategories (see Table IV and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
TABLE IV: Taxonomy of solutions of issues in microservices systems
@@
{tcolorbox}
[sharp corners, boxrule=0.1mm,] Key Findings of RQ3: We found that not all the issue discussions provide the information about their solutions, and finally we identified 1,899 issue discussions containing information about the solutions with 36 solutions mentioned by the interviewees and 1 solution mentioned by the survey participants, which are 1,936 solutions in total. The solution taxonomy consists of 8 categories, 32 subcategories, and 177 types of solutions. The majority of solutions are related to Fix Artifacts (54.54%), Add Artifacts (18.59%), and Modify Artifacts (10.95%).

### v3

@@ Abstract
keywords:
[…]
@@ 1 Introduction
Motivating Example: We now contextualize the issues, causes, and solutions based on an example illustrated in Figure 1. The example is taken from the Spinnaker project, an open-source microservices project hosted on GitHub (see Table 1), and annotated with numbering to represent a sequence among the reported issue, its cause(s), and the solution(s) to resolve the issue. Figure 1 shows auxiliary information about the Spinnaker project, such as project description, stars, and contributors. As shown in the example, a contributor, typically a microservices developer, writes code and may provide additional details of the code through comments. Once the code is compiled, the contributor reports permission denied issue highlighting “pipeline save with Admin account fails with permission denied”. As the next step, the same or other contributor highlights the cause for such issue as “Spinnaker user does not have access to the service account”. As the last step, an individual or a community of developers provides a solution such as “Allow the admin users to save the accounts” to resolve the issue. We are only interested in analyzing issues that have been marked as closed to ensure that a solution to resolve the issue exists. As shown in Table 1, the Spinnaker project has 4,595 closed issues and 121 contributors.
@@ 2.1 Research Questions
RQ1: What issues do occur in the development of microservice systems?
@@
RQ2: What are the causes of issues that occur in microservices systems?
Rationale: The aim of RQ2 is to investigate and classify the root causes behind the issues identified in RQ1 and map causes to issues. The answer to RQ2 helps practitioners avoid common issues in microservices systems.
RQ3: What solutions are proposed to fix issues that occur in microservices systems?
Rationale: The aim of RQ3 is to identify the solutions for the issues according to their causes and to develop the taxonomy of solutions. The answer to RQ3 helps to understand the fixing strategies for addressing microservices issues.
RQ4: What are the practitioners’ perspectives on the taxonomies of the identified issues, causes, and solutions in microservices systems?
Rationale: The taxonomies of issues, causes, and solutions developed in RQ1, RQ2, and RQ3 are based on 15 open-source microservices systems and interviewing 15 practitioners. RQ4 aims to evaluate the taxonomies of issues, causes, and solutions built in RQ1, RQ2, and RQ3 by conducting a relatively large-scale online survey.
@@ 2.2.1 Step A – Identify & Select MSA-based OSS Projects
We contacted a total of 426 contributors, with 39 of them responded to our query. Based on their confirmation, we selected 15 MSA-based OSS projects (see Table 1). These 15 selected projects, based on the available data at the time of project selection, represent the highest number of issues, making them highly relevant to our study.
Table 1: Identified open-source microservices systems
@@
Spinnaker 4595 121 1.2K 8.5K
Cortex 1120 226 681 4.7K
Jaeger 995 239 1.9K 15.8K
eShopOnContainers 986 157 8.9K 20.7K
Goa 930 97 485 4.7K
Light-4j 584 37 588 3.4K
Moleculer 473 102 497 5.1K
Microservices-demo 287 55 2.1K 3.1K
Cliquet 207 19 20 65
Deep-framework 174 12 75 537
Scalecube 130 20 90 547
Lelylan 123 7 93 1.5K
Open-loyalty 175 14 80 300
Spring PetClinic 69 32 1.4K 1.1K
Pitstop 39 15 490 890
@@ 2.2.2 Step B – Synthesize Issues, Causes, and Solutions
Raw Data Collection: We chose 15 microservices projects (see Table 1) as the source for building the dataset to answer our RQs. These 15 projects were chosen because they are significantly larger than other microservices projects hosted on GitHub. Hence, it is highly likely that their contributors had more discussions about the type of issues, causes, and solutions in issue tracking systems. The discussions relating to a software system can usually be captured in issue tracking systems [24]. We initially extracted 10,222 issue titles, issue links, issue opening and closing dates, and the number of contributors for each issue through our customized Python script (see the Raw Data sheet in [19]). We stored this information in MySQL and exported it into MS Excel sheets for further processing. We extracted only closed issues because it could increase the chances of answering all our RQs (e.g., solutions).
@@ 2.3.3 Data Analysis
Table 2: Overview of the interviewees and their demographic information
@@
P1 Software Architect, Developer Python, Java, Node.JS E-commerce, Healthcare 14 Years 6 Years Sweden
P2 Developer Java with Spark E-commerce 9 Years 5 Years USA
P3 Software Architect Python, Go, Java E-commerce 13 Years 6 Years UK
P4 Software Architect, Developer Python, Java Educational ERP 7 Years 3 Years Pakistan
P5 Software Architect Python, Java Internet of Things 12 Years 7 Years Canada
P6 Software Architect, Developer Java, Node.JS, Python Healthcare 15 Years 8 Years Australia
P7 Software Engineer C#.Net E-commerce, Banking 10 Years 4 Years Canada
P8 DevOps Consultant Kotlin, Python Network applications 9 Years 6 Years Canada
P9 Software Architect Java Education, Healthcare 5 Years 4 Years Chile
P10 Application Developer, Architect Java Financial (Insurance) 10 Years 8 Years Sweden
P11 DevOps Consultant Java, Kotlin,Python Telecommunication 6 Years 3 Years Norway
P12 Application Developer Angular, Puthon Manufacturing 8 Years 4 Years China
P13 Principal Consultant Swift Transportation 9 Years 5 years USA
P14 Azure Technical Engineer JavaScript, Golang Embedded systems 12 Years 4 Years Pakistan
P15 Software Architect Ruby, UML Payment applications 16 Years 7 Years India
@@ 3 Results
This section presents the analyzed results of this study, addressing the four RQs outlined in Section 2.1. The analyzed results are further organized as categories (e.g., Technical Debt), subcategories (e.g., Code Debt), and types (e.g., code smell). We present categories in boldface, subcategories in italic, and types in small capitals. The relevant examples are provided as quoted messages along with their issue ID numbers to facilitate the traceability to our dataset (see the Initial Codes sheet in [19]). We report the types of issues in Section 3.1 (see Figure 4), the types of causes in Section 3.2 (see Table 3), the types of solutions in Section 3.3 (see Table 4), and practitioners’ perspective on these taxonomies in Section 3.4 (see Table 5).
@@ 3.1 Types of Issues (RQ1)
Key Findings of RQ1: We identified 2,641 instances of issues by mining developer discussions in 15 open-source microservices systems with 48 instances of issues mentioned by the interviewees and 9 instances of issues mentioned by the survey participants, which are 2,698 issues in total. The issue taxonomy consists of 19 categories, 54 subcategories, and 402 types, indicating the diversity of the issues in microservices systems. The majority of issues are related to Technical Debt (25.46%), CI/CD (11.60%), and Exception Handling (8.45%).
@@ 3.2 Causes of Issues (RQ2)
The taxonomy of causes of microservices issues is provided in Table 3. It is worth mentioning that not all the issue discussions provide the information about their causes. Therefore, we identified 2,225 issue discussions containing information about the causes. The taxonomy of causes is derived by mining developer discussions (i.e., 2,225 instances of causes), conducting practitioner interviews (i.e., 31 instances of causes, see Section 2.3.3), and conducting a survey (i.e., 11 instances of causes, see Section 3.4). Hence, we got a total of 2,267 instances of causes. We identified a total of 228 types of causes that can be classified into 8 categories and 26 subcategories. Due to space limitations, we only list the top two types of causes for each subcategory in Table 3. The detail of the types of causes can be found in the dataset [19]. The results show that General Programming Error (860 out of 2267), Missing Features and Artifacts (386 out of 2267), and Invalid Configuration & Communication Problems (382 out of 2267) are the top three categories of causes. Each cause category is briefly discussed below.
1. General Programming Error (GPE) (860/2267, 37.93%): This category captures the causes that are based on a broad range of errors occurred in different phases of microservices system development, such as coding (e.g., syntax errors), testing (e.g., incorrect test cases), and maintenance (e.g., wrong examples in documentation) phases. We identified and classified 78 types of GPE causes in six subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
2. Missing Features and Artifacts (MFA) (386/2267, 17.02%): This category represents the causes behind microservices systems issues that occur due to missing required features, packages, files, variables, and documentation and tool support. We identified and classified 27 types of MFA causes in four subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 29 types of causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 28 types of causes in five subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 17 types of SD&IA causes in 2 subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 21 types of PSM causes in 3 subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
7. Insufficient Resources (IR) (93/2267, 4.10% ): Microservices systems are at risk of delivering the required outcome without sufficient resources. We identified and classified 4 types of IR causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
8. Fragile Code (FC) (24/2267, 1.05%): Fragile code refers to code that is difficult to change, and a minor modification in fragile code may break the service or module. We identified and classified 6 types of FC causes in two subcategories (see Table 3 and the Cause Taxonomy sheet in [19]). Each of them is briefly described below.
@@
Table 3: Taxonomy of causes of issues in microservices systems
@@
Key Findings of RQ2: We found that not all the issue discussions provide the information about their causes, and finally we identified 2,225 issue discussions containing information about the causes with 31 causes mentioned by the interviewees and 11 causes indicated by the survey participants, which are 2,267 causes instances in total. The cause taxonomy of microservices issues consists of 8 categories, 26 subcategories, and 228 types. The majority of causes are related to General Programming Errors (37.93%), Missing Features and Artifacts (17.02%), and Invalid Configuration and Communication (16.85%).
@@ 3.3 Solutions of Issues (RQ3)
The taxonomy of solutions for microservices issues is provided in Table 4. It is worth mentioning that not all the issue discussions provide the information about their solutions. Therefore, we identified 1,899 issue discussions containing the information about the solutions. The taxonomy of solutions is derived by mining developer discussions (i.e., 1,899 solutions), conducting practitioner interviews (i.e., 36 solutions, see Section 2.2.2), and conducting a survey (i.e., one instance of solution, see Section 3.4). Hence, we got a total of 1,936 solutions. We identified a total of 196 types of solutions that can be classified in 8 categories and 35 subcategories. Due to space limitations, we only list the top two types of solutions for each subcategory in Table 4. The detail of the types of solutions can be found in the dataset [19]. The results show that Fix Artifacts (1056 out of 1936), Add Artifacts (360 out of 1936), and Modify Artifacts (210 out of 1936) are the top three categories of solutions. Each solution category is briefly discussed below.
1. Fix Artifacts (1056/1936, 54.54%): During the analysis of developer discussions about microservices issues, we identified that most developers did not explicitly mention any solution to the problems. They fixed the issue in the local repository and send the fixed code (e.g., through a pull request) to the maintainer of the public repository. In this case, from the developer discussions, we could not find exactly what they added, removed, or modified in the project to fix a specific issue. Therefore, we named this category Fix Artifacts. We identified and classified 25 types of fix artifacts solutions in 4 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 33 types of solutions in 10 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
3. Modify Artifacts (212/1936, 10.95%): Besides adding new artifacts to the existing system to address the microservices issues, we also identified a large number of solutions in which the developers explicitly mentioned how they modified modules, services, packages, APIs, scripts, methods, objects, data types, identifiers, databases, and documentation to address the microservices issues. We identified and classified 31 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
4. Remove Artifacts (39/1936, 2.01%): This category covers the solutions in which artifacts are removed to address several types of microservices issues. We identified and classified 14 types of solutions in 5 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
5. Manage Infrastructure (160/1936, 8.26%): This category captures the solutions that are based on efficient resource utilization for addressing the microservices issues. We identified and classified 20 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
6. Manage Configuration and Execution (50/1936, 2.58%): Managing configuration and execution enables developers to track changes in microservices systems and their consuming applications over time, for example, the ability to track the version history of configuration changes for multiple instances of microservices systems. We identified and classified 13 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
We identified and classified 21 types of solutions in two subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
8. Import/Export Artifacts (12/1936, 0.61%): We found several issues that can be fixed by importing and exporting various artifacts. We identified and classified 3 types of solutions in 2 subcategories (see Table 4 and the Solution Taxonomy sheet in [19]). Each of them is briefly described below.
@@
Table 4: Taxonomy of solutions of issues in microservices systems
@@
Key Findings of RQ3: We found that not all the issue discussions provide the information about their solutions, and finally we identified 1,899 issue discussions containing information about the solutions with 36 solutions mentioned by the interviewees and 1 solution mentioned by the survey participants, which are 1,936 solutions in total. The solution taxonomy consists of 8 categories, 32 subcategories, and 177 types of solutions. The majority of solutions are related to Fix Artifacts (54.54%), Add Artifacts (18.59%), and Modify Artifacts (10.95%).